### v14

# D4C: Improving Negative Example Quality to Enhance Machine Abstract Reasoning Ability

Abstract
This paper is dedicated to addressing the challenge of enhancing the abstract reasoning capabilities of deep learning algorithms, particularly for tasks involving complex human concepts. We introduce Lico-Net, a novel reasoning engine grounded in fuzzy system theory, which encodes the logical structure of Raven’s Progressive Matrices (RPM) problems into fuzzy representations. Lico-Net excels in solving RPM tasks. Furthermore, we propose Lico-Net-Bongard, a tailored version of Lico-Net for the Bongard-Logo problem, which also achieves high reasoning accuracy through fuzzy representations. However, we observe a mismatch between the way deep learning algorithms and humans induce reasoning concepts, primarily attributed to the inadequate quality of negative samples. Improper configuration of negative samples can convey erroneous conceptual information to deep learning algorithms, thereby distorting their learning objectives. To address this issue, we propose two novel approaches: first, treating different sample points within reasoning problems as mutual negative samples to alter the existing negative sample structure in the data; second, designing a negative sample generator based on a step-wise linear attention mechanism to produce high-quality negative samples. Experimental results demonstrate that these methods significantly improve the performance of Lico-Net (-Bongard) and other baseline models on the RPM and Bongard-Logo datasets, as well as in the domain of foundational vision model processing, particularly when addressing the NICO dataset’s distribution shift problem. Our findings emphasize the importance of improving negative sample quality for enhancing the abstract reasoning capabilities of deep learning algorithms and suggest that fuzzy systems represent a promising direction for future research in this field.
Index Terms:
Abstract Reasoning, Raven’s Progressive Matrices, Bongard-logo, NICO, Non-Independent and Identically Distributed, Adversarial Learning, Fuzzy System.
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
Deep learning, as an artificial intelligence technology that mimics the human brain’s operating mechanism, has shone brightly in numerous fields. In generative tasks, it has demonstrated extraordinary talent in image generation [9], text creation, and video production through cutting-edge technologies such as Generative Adversarial Networks (GANs) [7], Variational Autoencoders (VAEs) [8], autoregressive models, and Transformer [4]. Meanwhile, deep neural networks have also achieved remarkable success in various domains, including computer vision [1, 2, 3], natural language processing [4, 5, 6], generative model construction [7, 8, 9], visual question answering systems [10, 11], and visual abstract reasoning [16, 12, 13]. However, deep learning faces challenges in discriminative tasks, which require models to accurately classify or regress based on input data, such as image classification, text categorization, and speech recognition. Despite some achievements in these tasks, deep learning still struggles to deeply understand the human concepts embedded in the data.
This dilemma may stem from deep learning models’ overreliance on superficial statistical patterns in training data, neglecting the deeper structures and logical relationships that underlie human concepts. Therefore, despite the widespread success of deep learning, we still have a long way to go in truly understanding and simulating complex and vague human concepts. This limitation is particularly evident when dealing with data that contains highly complex and ambiguous human concepts.
Given this, our research will focus on discriminative problems involving complex and deep categorical concepts. We will delve into the root causes of deep learning models’ poor performance on such issues and actively explore effective solutions, aiming to drive new breakthroughs in deep learning’s ability to understand and simulate human concepts.
II Problem description
Traditional visual discriminative tasks evaluate the generalization and feature recognition capabilities of deep learning algorithms in terms of pixel arrangements within image data. However, in this section, we introduce several advanced visual discriminative tasks that deviate from the norm. These tasks effectively gauge the ability of deep learning algorithms to learn complex and abstract patterns underlying the visible features. This paper presents a series of fuzzy systems that have excelled in these challenging tasks, highlighting the significance of fuzzy systems as a crucial branch of artificial intelligence.
II-A Raven’s Progressive Matrices
RPM problems, or Raven’s Progressive Matrices, are an authoritative benchmark for assessing the advanced discriminative capabilities of deep learning algorithms due to their requirement for abstract reasoning, pattern recognition, and problem-solving skills. As a high-level, non-verbal task that demands the ability to generalize and transfer knowledge, RPM challenges algorithms to demonstrate cognitive functions akin to human intelligence. Its standardized nature and strong correlation with human intelligence measures make it a relevant and consistent test for evaluating the sophistication of machine learning models, emphasizing its authority in gauging the intelligent level of deep learning algorithms.
II-A1 RAVEN and PGM
The RAVEN database is a collection of Ravens Progressive Matrices problems, each consisting of 16 images. Eight of these images, known as the “statement,” are the main challenge, while the other eight form a pool for selection. The goal is to choose the right images to complete a $3\times 3$ matrix with a progressive geometric pattern, following an abstract ”rule.” The left side of Figure 1 shows an example of RAVEN. Similarly, PGM provides eight images for the stem and answers, but its ”rule” applies to both rows and columns. The right side of Figure 1 gives a typical example of PGM.
Figure 1: RAVEN and PGM case
II-B Bongard problem
The Bongard problem[12, 13] is a challenging abstract visual reasoning task that assesses pattern recognition by presenting images following one of two contrasting rules. It requires a limited number of training examples and demands the ability to abstract and learn from minimal data, posing a significant challenge for both humans and machine learning algorithms. Despite being a well-recognized test in AI and cognitive science, there is limited literature on automatically solving Bongard problems, which involve not just pattern recognition but also higher cognitive functions like abstract reasoning and concept learning, where current machine learning methods significantly trail human capabilities.
II-B1 Bongrad-logo
The Bongard-Logo[13] problem extends the Bongard problem, serving as a benchmark for human-level concept learning and reasoning. It aims to capture three key aspects of human cognition: context-dependent perception, analogical perception, and perception with limited samples but vast vocabulary. Described using action-oriented LOGO language, Bongard-Logo problems can generate thousands of instances via program-guided techniques. They challenge standard image classification methods’ objectivism and metaphysical realism, necessitating a new, human-like perception. Experimental results show that advanced deep learning methods fall short of human performance, indicating they miss core human cognitive characteristics. Bongard-Logo problems guide research towards a general visual reasoning architecture, aiming to surpass this benchmark.
Figure 2 illustrates a representative Bongard-logo case, where every Bongard case consists of two sets of images: Set A (primary) and Set B (auxiliary). Set A contains six images, with geometric entities in each image adhering to a particular set of rules, while Set B encompasses six images that do not conform to the rules observed in Set A. The task entails assessing whether the images in the test set align with the rules manifested in Set A. The level of difficulty varies depending on the case’s structural complexity.
Figure 2: Bongard-Logo case
Bongard-logo problems can be categorized into three parts based on concept domains: 1) Free-form problems (FF), featuring shapes made of random action strokes; 2) Basic shape problems (BA), recognizing pattern of shape combinations; 3) High-level concept problems (HD), assessing abstract concept reasoning, like convexity and symmetry.
Both RPM problems and Bongard-Logo problems are challenging for humans with extensive prior knowledge, and deep learning algorithms face even higher cognitive demands in the absence of prior knowledge. Deep learning algorithms that stand out in such environments typically bring methodological breakthroughs.
II-C Out-of-distribution problems in image backgrounds
Out-of-distribution (O.O.D.) problems in image backgrounds involve the challenge of identifying images with backgrounds that differ significantly from the training data, which is crucial for maintaining the reliability of machine learning models in real-world applications. These problems are particularly relevant in high-stakes fields such as medical diagnosis and autonomous driving, where misclassification can have severe consequences. Addressing O.O.D. issues often relies on learning representations that distinguish in-distribution data from O.O.D. samples, with techniques like Masked Image Modeling showing promise in improving detection performance by capturing the intrinsic distributions of the training data.
II-C1 NICO database
The NICO database[19] is specifically designed to address the issues of Non-I.I.D. (Non-Independent and Identically Distributed) or Out-of-Distribution (O.O.D.) image classification problems. It simulates scenarios where the test distribution may deviate arbitrarily from the training distribution, violating the traditional I.I.D. assumption of most machine learning methods. The core idea of the NICO dataset is to label images with a main concept (e.g., dog) and context (e.g., grass), and by adjusting the proportion of different contexts in training and test data, it flexibly controls the degree of distribution shift and enables research under various Non-I.I.D. settings.
The NICO dataset includes two main categories: NICO-animals and NICO-vehicles, with 10 animal categories and 9 vehicle categories, each having 9 or 10 contexts. Some examples of the NICO-animal and NICO-vehicle are shown in Figure 3.
Figure 3: The figure shows the example of NICO-animal and NICO-vehicle.
This paper is dedicated to exploring a deep learning approach that does not rely on pre-training or contextual prompts.
III Related work
III-A RPM solver
Image reasoning models like CoPINet[22], LEN+teacher[21], and DCNet[23] emphasize learning disparities and rules, while NCD[25], SCL[26], and SAVIR-T[27], along with neural symbolism systems (PrAE, NVSA, and ALANS[28, 29, 30]), enhance interpretability and accuracy. RS-CNN and RS-TRAN[31] have particularly excelled at solving RPM problems. Triple-CFN[37] distinguishes itself by implicitly extracting and indexing concept and reasoning unit information, leading to significant reasoning accuracy. CRAB[32], based on Bayesian modeling, has made significant advancements with a customized conservatory for the RAVEN database, despite not tackling core RAVEN challenges, and the scientific community eagerly awaits the broader impact of such innovative methods.
III-B Bongard solver
Addressing Bongard problems, approaches include language feature models, CNNs, and synthetic data. Depweg et al.[12] used Bayesian inference with formal language but faced scalability issues. Kharagorgiev and Yun[35] focused on feature extraction, and Nie et al.[13] experimented with CNNs and meta-learning with limited success. PMoC[36] offers a probabilistic approach to assess group sample likelihoods, proving effective. Triple-CFN[37] stands out for its consistent effectiveness on both Bongard-Logo and RPM problems.
III-C NICO solver
Stable learning [33] addresses the O.o.D. problem by sample reweighting to correct domain distribution shifts and reduce feature correlations, enhancing model performance on unseen test sets. DecAug [34] tackles the two-dimensional O.o.D. problem in NICO data by introducing domain labels, splitting the deep learning model into branches for domain and object learning, with perpendicular parameter learning directions to ensure information decoupling, albeit increasing the workload and categorization difficulty. SimCLR [43] improves representation learning through data augmentation, a learnable nonlinear projector, and the InfoNCE loss function. VICReg [42] introduces three regularization terms on embedded variance for self-monitored image representation learning.
III-D Distribution distance measurement method
III-D1 Wasserstein distance
The Wasserstein distance [38], a metric for distribution dissimilarity, is given by:
$W(P,Q)=\mathbf{inf}\left(\sum|X-Y|\right)$
where $P$ and $Q$ are probability distributions, and $X$ and $Y$ are random variables from $P$ and $Q$, respectively. The distance $|X-Y|$ can be Euclidean or another metric. It quantifies the minimum cost of transforming one distribution into another and is widely used in generative models to assess the difference between generated and real images.
III-D2 Sinkhorn distance
Sinkhorn distance[39] is a metric based on optimal transport, which approximates the Wasserstein distance through the introduction of entropy regularization, leading to increased computational efficiency.
IV Methodology
In this study, we emphasize that deep learning algorithms should focus on learning the fuzzy representations of complex and tightly intertwined human concepts when addressing such issues. By leveraging these fuzzy representations to solve problems, the performance of the algorithms can be significantly enhanced. However, we have observed that, despite the performance boost brought by the introduction of fuzzy representations, there exists a notable discrepancy between the way deep learning algorithms summarize concepts and logical patterns and that of humans. Upon thorough analysis, we found that this mismatch is not attributable to the fuzzy representation technique itself but is a widespread issue in traditional deep learning algorithms. Ultimately, one of the core reasons lies in the inadequate configuration of negative samples in the data. Improper negative sample configuration can convey erroneous conceptual information to deep learning algorithms, thereby distorting their learning objectives.
Therefore, this study posits that, in addition to striving to develop deep learning algorithms capable of effectively learning from low-quality negative samples, improving the quality of negative samples is also a crucial strategy.
V Discriminator of Concepts (DC)
In this section, we introduce Lico-Net, a novel deep neural network-based inference engine designed to efficiently solve RPM (Raven’s Progressive Matrices) problems. By leveraging advanced techniques, Lico-Net fuzzifies the concepts and reasoning logic embedded in RPM problems, achieving excellent performance. Subsequently, we present Loco-Net-Bongard, a deep learning inference engine tailored for solving Bongard-Logo problems, which also attains outstanding reasoning accuracy through explicit concept fuzzification. Furthermore, preliminary experiments on Lico-Net (and its Bongard counterpart) reveal a phenomenon of concept mismatch, highlighting the models’ limitations in learning implicit concepts. Additionally, using the NICO dataset, we further demonstrate how shifts in image background distributions can obscure deep learning models’ ability to learn implicit concepts.
V-A A novel baseline, Lico-Net, for RPM problems
The RPM problem requires participants to select answers from an option pool in order to complete a $3\times 3$ progressive matrix of images. If the completed matrix adheres to a specific image progressive pattern that can be inferred from the incomplete matrix (problem statement), it is considered a correct answer selection[41]. Therefore, a successful RPM discriminator is expected to be able to identify the rationality of the progression patterns within a complete $3\times 3$ progressive matrix.
To address this challenge and enhance the ability to discern underlying patterns, we have devised a novel method. We propose Lico-Net (link-conception net), a network specifically designed for RPM (Raven’s Progressive Matrices) reasoning engines. The core function of Lico-Net is to evaluate the consistency of progressive patterns within a 3×3 RPM matrix and quantify this consistency with a score. By sequentially applying each option provided in an RPM case to complete the matrix, we can utilize Lico-Net to assess the resulting progressive consistency and assign scores to each option accordingly. This approach enables us to efficiently reason through and solve RPM problems. The following is an overview of the detailed design of Lico-Net.
The macro architecture of the Lico-Net we designed is presented in Figure 4. As can be seen from the figure, Lico-Net is divided into four core modules: the Perception Module, the Inference Module, the Concept Fuzzification Module, and the Concept Consistency Evaluation Module. Figure 4 intuitively illustrates our concept; we plan to use the Perception Module to convert all images in the RPM matrix into corresponding representations. Subsequently, the Inference Module and the Concept Fuzzification Module will explore the concepts implicit in the representation matrices with missing entries and depict these concepts using fuzzy representations. Finally, through the Concept Consistency Measurement Module, we will accurately assess the consistency of these concepts. The specific design of the four modules is as follows:
Figure 4: The figure shows the framework of Lico-Net.
V-A1 Perceptron module of Lico-Net
Aligned with the consensus among RPM solvers on the importance of multi-viewpoint feature extraction [31, 26, 27, 37], Lico-Net employs the Vision Transformer (ViT) for feature extraction, retaining the entire output vectors, which are attention results, as multi-viewpoint features. We denote the number of viewpoints as $L$. Lico-Net then equally processes each viewpoint of the extracted feature. The perceptron of Lico-Net is illustrated in Figure 5.
Figure 5: Perceptron of Lico-Net
V-A2 Inference module and Concept Fuzzification Module of Lico-Net
The inference module of Lico-Net processes the representation matrices generated by the perception module for each viewpoint equally and in parallel, resulting in $L$ sets of concepts represented in the form of distributions. During the concept extraction process, the inference module enumerates all possible ways in which the representation matrix can have missing entries, then extracts the concepts implied by each incomplete matrix and represents these concepts as distributions. Ultimately, a distribution distance measurement function is utilized to calculate the distances between concepts within each set, serving as a measure of conceptual consistency. We have illustrated in Figure 6 how the inference module handles a specific missing entry configuration in the representation matrix and transforms it into a concept represented by a distribution.
Figure 6: The process by which the inference module handles the third incomplete representation matrix to extract concepts.
On the left side of the figure 6, it is easy to see that in the Inference module, the MLP with a bottleneck structure extracts the remaining row and column information from the incomplete matrix, which is then fed into a standard Transformer-Encoder to extract the concepts present in the incomplete matrix. The figure shows that we use the Transformer-Encoder to extract 5 feature vectors from 4 row and column informations to represent the concepts, a kind of asymmetric token transformation that relies on learnable vector embeddings, which has been used in ViT. On the right side of the figure, the Concept Fuzzification Module is depicted, illustrating the method of converting vector-form concepts into fuzzy concepts. This involves using a fuzzification head to double the dimensionality of the vectors, then splitting them and treating each part as the mean ($\mu$) and the logarithm of the variance ($\log\sigma^{2}$) of a Gaussian distribution. Subsequently, the reparameterization trick is employed to obtain fuzzy representations of these concepts based on the encoded mean ($\mu$) and variance ($\sigma^{2}$). In this design, both the concept’s baseline ($\mu$) and the degree of fuzziness ($\sigma^{2}$) are learned autonomously. The design of the fuzzification head is shown in Figure 7.
Figure 7: The figure shows that the structure of fuzzification head.
V-A3 Concept Consistency Measurement Module
The inference module is capable of transforming incomplete representation matrices into sampled results of Gaussian Mixture Models (GMMs) with five components, treating these samples as the fuzzy representations of inherent concepts within the incomplete matrices. A complete $3\times 3$ matrix can give rise to nine distinct incomplete matrices; hence, we can extract nine sets of concepts from a single complete matrix, obtaining their fuzzy representations. Given that these fuzzy representations of concepts follow a mixture of Gaussian distributions, the task of the Concept Consistency Measurement Module is to calculate the distances between these GMMs based on the Sinkhorn distance [39], and to use this distance to reflect the consistency score of the concepts within a complete $3\times 3$ matrix.
V-A4 Loss function for Lico-Net
By employing cross-entropy loss and constraining this consistency score according to the logic of the problem, we can end-to-end optimize the three modules of Lico-Net.
V-A5 Preliminary experiments in Lico-Net
After clarifying the structure of Lico-Net and its training loss function, we conducted preliminary experiments across all sub-datasets of RAVEN. Rather than focusing solely on the inference accuracy of Lico-Net, we placed greater emphasis on how Lico-Net learns the embedded concepts within the RAVEN dataset. Since Lico-Net can extract nine sets of concepts from a single complete $3\times 3$ matrix, it is worth delving into its concept extractor’s performance across the entire dataset.
Therefore, we organized this special experiment. Taking the “Distribution 9” configuration in RAVEN as an example, 20 problem matrices with identical rules were generated for this configuration. Subsequently, we assessed the distances between concepts within and between these 20 problem matrices. As shown in Figure 8, despite the identical rules for these 20 problem matrices, it is observed from Lico-Net’s perspective that only the concepts within the same matrix are consistently matched. In other words, Lico-Net’s induction of concepts or patterns is far from human-like. Whether this pattern mismatch is a negative effect brought by concept fuzzification is worth discussing.
Figure 8: The figure shows that the distances between concepts within and between these 20 problem matrices that share the same rule.
V-B A novel baseline, Lico-Net-Bongard, for Bongard-Logo
The designers of Bongard-logo drew inspiration from high-dimensional human concepts and knowledge, carefully categorizing the Bongard-logo database into three distinct problem types: FF, BA, and HD. It is noteworthy that during the testing process, the HD concept is further decomposed into NV and CM. Bongard-Logo requires participants to analyze the clustering logic between the primary and auxiliary sets outlined in the problem statement, and to correctly categorize two given images into these sets. For this type of problem, our approach, which involves extracting concepts from images as fuzzy representations, is likely to be effectively applied. By utilizing a deep neural network to extract and represent concepts from images as fuzzy representations, we can naturally impose constraints on the distances between these concepts to align with the problem-solving logic and clustering logic of Bongard-Logo. This ensures that concepts within the primary set are more closely related, while concepts within the auxiliary set are distinct from those in the primary set.
Based on our methodology, we have designed Lico-Net-Bongard.
V-B1 The structure of Lico-Net-Bongard
For a clearer exposition of the Lico-Net-Bongard architecture, Figure 9 has been constructed. In Figure 9 (a), we detail the annotation of the 14 images within a single Bongard-Logo problem, which is instrumental in defining the loss function for training Lico-Net-Bongard. Figure 9 (b) illustrates the process by which Lico-Net-Bongard extracts concepts from individual images in Bongard-Logo. The process is as follows: after utilizing ResNet18 to convert Bongard-Logo images ($x_{i}$) into feature maps ($h_{i}$), Lico-Net-Bongard scatters these feature maps along the receptive field into vector groups ($h_{ij}$ where $j$ denotes the receptive field index and $m$ represents the size of the receptive field), which are then processed by a standard Transformer-Encoder to obtain the multi-viewpoint representations ($h^{\prime}_{ij}$) of the image. Thereafter, using a fuzzification head, these multi-viewpoint representations are individually transformed into distributions ($z_{ij}$). Additionally, reparameterization techniques are employed to compute the fuzzy representations of the concepts from the transformed distributions.
Figure 9: The structure of Lico-Net-Bongard
V-B2 Loss function for Lico-Net-Bongard
After Lico-Net-Bongard processes all 14 images ($\{x_{i}|\,i\in[1,14]\}$) in a Bongard-Logo case to form concepts ($\{z_{ij}|\,i\in[1,14],j\in[1,m]\}$), we can optimize Lico-Net-Bongard by constraining the distances between these concepts. The specific loss function design can be expressed as:
${\ell_{\mathbf{Lico}}}$
$=\sum_{i=1}^{6}\sum_{\tilde{i}=i+1}^{7}-\log\frac{{{e^{\frac{\sum\nolimits_{j=1}^{m}D({z_{ij}},{z_{\tilde{i}j}})}{m}}}}}{{{e^{\frac{\sum\nolimits_{j=1}^{m}D({z_{ij}},{z_{\tilde{i}j}})}{m}}}+\sum\nolimits_{k=8}^{14}{{e^{\frac{\sum\nolimits_{j=1}^{m}D({z_{ij}},{z_{kj}})}{m}}}}}}$
$+\sum_{i=1}^{6}\sum_{\tilde{i}=i+1}^{7}-\log\frac{{{e^{\frac{\sum\nolimits_{j=1}^{m}D({z_{ij}},{z_{\tilde{i}j}})}{m}}}}}{{{e^{\frac{\sum\nolimits_{j=1}^{m}D({z_{ij}},{z_{\tilde{i}j}})}{m}}}+\sum\nolimits_{k=8}^{14}{{e^{\frac{\sum\nolimits_{j=1}^{m}D({z_{\tilde{i}j}},{z_{kj}})}{m}}}}}}$ (1)
The expression for $\ell_{\mathbf{Lico}}$ is straightforward: it calculates the distances between each pair of concepts within the set $\{z_{ij}\}$ and constrains these distances using cross-entropy loss. Here, $D(\alpha,\beta)$ represents the function that calculates the Sinkhorn distance between two distributions $\alpha$ and $\beta$. Similarly, we employ the reparameterization method to sample from $z_{ij}$, with the number of samples set to 10.
V-B3 Preliminary experiments in Lico-Net-Bongard
With the structure of Lico-Net-Bongard and its training loss function clearly defined, we conducted preliminary experiments on the Bongard-Logo dataset. We carried out two sets of experiments, one where the four types of concepts from the Bongard-Logo database were mixed together, and another where they were separated, to observe the performance of Lico-Net-Bongard. The experimental results are presented in Table I.
TABLE I: Reasoning Accuracies of Lico-Net-Bongard on Bongard-logo.
Test Accuracy(%)
Model and Data Set FF BA CM NV
Lico-Net-Bongard on Bongard-logo 91.0 98.3 76.8 77.0
Lico-Net-Bongard on Separated Bongard-logo 98.8 99.3 75.5 73.4
As observed in Table I, when the concepts in the Bongard-Logo dataset are artificially separated, Lico-Net-Bongard’s performance improvement in BA and FF, which are independent and identically distributed concepts, and the performance decline in NV and CM, which are generalization concepts, both reflect the situation where Lico-Net-Bongard confuses these four concepts.
V-C The baseline on the NICO database.
The NICO database presents a classic visual discrimination task, but the data within it exhibit a shift between the distribution of background pixels and the distribution of foreground subjects. In NICO, the “domain” refers to the background element, which can undergo two types of shifts: Diversity Shift and Correlation Shift. Diversity Shift occurs when the test set contains new, unseen background styles. For example, a model trained to recognize sheep may struggle if it encounters sheep in snow, a scenario not present in the training data. Correlation Shift, on the other hand, involves the relationship between background and foreground pixel distributions in the training set. If most sheep images in the training data have a grassland background, a model might mistakenly associate the label with the background rather than the foreground, leading to impaired performance.
V-C1 Preliminary experiments in NICO
From an experimental perspective based on the NICO dataset, we describe the issues of Diversity Shift and Correlation Shift.
Diversity Shift: We randomly select 8 domains as source domains and the remaining 2 as target domains. After training and testing basic deep learning models on all 10 domains to establish a benchmark, we train the model on source domains and test it on target domains. By comparing the results to the benchmark, we observe the impact of Diversity Shift. For a fair comparison, we maintain consistent data volumes in all training and testing sets. The experimental results in Table II highlight the significant effect of Diversity Shift on deep learning models.
TABLE II: Experiment of Diversity Shift Impact (8:2)
dataset model Domain mixing(10:10)1 Domain shift(8:2)2
NICO-Animal ResNet18 41.03$\pm$0.1028.56$\pm$0.30%
ResNet50 54.55$\pm$0.1539.55$\pm$0.25%
Vgg16 62.49$\pm$0.2148.66$\pm$0.50%
Vgg19 60.22$\pm$0.1346.11$\pm$0.09%
NICO-Vehicle ResNet18 53.45$\pm$0.1242.29$\pm$0.08%
ResNet50 70.20$\pm$1.1855.16$\pm$0.35%
Vgg16 74.13$\pm$1.3358.24$\pm$0.40%
Vgg19 74.59$\pm$1.2657.60$\pm$0.80%
  • 1
    Separate training set and test set from mixed 10 domians.
  • 2
    The training set is from the source domain, and the test set is from the target domain.
Correlation Shift: In our experiment, we exacerbate this shift by randomly downsampling the training set. Specifically, we randomly select one of the 8 source domains as the primary and reduce the sample sizes of the remaining 7 to 1/5 of the primary. We then train on the reduced source domains and test on the unchanged target domain. Table III shows Vgg16 results on the NICO dataset with more pronounced correlation shift.
TABLE III: Vgg16 on nico 8:2 with Correlation shift
dataset model original (8:2)1 Stronger domain shift(8:2)2
NICO-Animal vgg16 48.66$\pm$0.5037.83$\pm$0.12%
NICO-vehicle vgg16 58.24$\pm$0.4041.85$\pm$0.25%
  • 1
    Do not change the data volume proportion of the source domain.
  • 2
    The data volume of the secondary domain in the source domain is reduced by $1/5$ of the primary domain.
Since prevalent pre-training (e.g., ImageNet pre-trained models) covers most of the information needed for our experiments, none of the experiments on NICO adopt a pre-training scheme. We conduct each experiment 5 times and report the average of the 5 test accuracies.
V-D Discussion on Preliminary Experiments
In our preliminary experiments, there is a pervasive phenomenon that there exists a mismatch between the way deep learning models infer deep concepts and the way humans define these concepts.
In the Bongard-Logo problem, different concepts require Lico-Net-Bongard to store different image pixel encoding patterns. When these encoding patterns are mutually exclusive, the mixed concepts confuse Lico-Net-Bongard, making it uncertain which concept to rely on when solving problems during testing. This transforms four concepts that are clear and organized from a human perspective into four conflicting concepts from the perspective of Lico-Net-Bongard. As a result, the phenomenon of concept mismatch observed in the concept dissection experiment becomes understandable. However, the design of the correct answers in the RAVEN problem is indeed based on its inherent rules and problem-solving logic, and the annotations in NICO are also accurately aligned with the foreground subjects. Therefore, the question arises: why do both RAVEN and NICO exhibit a degree of conceptual mismatch?
VI Attribution Discussion of Mismatch in Deep-level Concepts (D2C)
Our preliminary experiments have revealed a significant finding: there exists a mismatch between the way our designed inference engine induces deep or implicit concepts in the data and the way humans interpret these concepts. This mismatch is one of the key factors limiting the performance of the inference engine. Therefore, delving into the causes of the concept mismatch and devising corresponding solutions will be an effective approach to enhancing the performance of the inference engine.
Upon analysis, one of the reasons for concept mismatch is attributed to the low quality of the negative examples set in the problems.
VI-1 RPM problem
Taking the RPM problem, most end-to-end RPM inference engine, including Lico-Net, follow a pattern: when the correct option is filled into the problem matrix, the inference engine returns a high probability value for this matrix, and conversely, returns a low probability value. This indicates that the essence of an RPM inference engine is to perform a conditional probability distribution over the options, conditioned on the problem matrix. The optimization of the inference engine can be seen as an adjustment of this conditional probability distribution. If we assume that the form of this conditional probability distribution, after determining the problem matrix, is a Gaussian distribution, then the optimization of this distribution can be regarded as adjusting the mean $\mu_{\theta}$ and variance $\sigma_{\theta}^{2}$ in the following formula:
$-\log(\text{Engine}(x_{\text{option}}|\,\text{matrix}))$
$=\frac{{(x_{\text{option}}-\mu_{\theta})^{2}}}{{2\sigma^{2}_{\theta}}}-\log(|\,\sigma_{\theta}|\,)-\frac{1}{2}\log(2\pi)$ (2)
Here, $x_{\text{option}}$ represents the options filled into the matrix, and $\mu_{\theta}$ and $\sigma_{\theta}^{2}$ represent the mean and standard deviation of the Gaussian distribution, respectively.
It can be observed that when positive samples (correct option) are substituted into the formula 2, we aim for the smallest possible result, which corresponds to minimizing the Euclidean distance between option $x_{\text{option}}$ and the mean of the correct solution $\mu_{\theta}$. Conversely, when negative samples (incorrect options) are substituted, we desire the largest possible result, equivalent to minimizing the standard deviation $\sigma_{\theta}^{2}$. It is well understood that the mean $\mu_{\theta}$ describes the center of a Gaussian distribution, while variance $\sigma_{\theta}^{2}$ describes the spread or dispersion. Consequently, it is not difficult to comprehend that the learning objective of end-to-end RPM inference engine like Lico-Net is a probability distribution delimited by positive and negative examples. In other words, the learning goal of these scoring-based end-to-end inference engine is not the Inference logic established by humans, but rather the distinction between positive and negative examples.
This indicates that when the quality of negative examples is not sufficiently high or not deceptive enough, it can lead to the engine learning concepts with relatively loose boundaries, which also causes the learned concepts to deviate significantly from the reasoning logic established by humans.
VI-2 Bongard-Logo problem
This issue has also impacted the process of solving and learning the Bongard-Logo problem. By examining Equation 1, it is evident that the primary learning focus of Lico-Net-Bongard remains on ensuring that the similarity among concepts extracted from positive examples is higher than that between positive and negative examples. In other words, its learning objective is to distinguish the differences between positive and negative examples, rather than directly grasping the four specific concepts defined by humans in Bongard-Logo. Consequently, due to the inadequate quality of the negative examples provided by Bongard-Logo, it communicates a set of conflicting concepts.
VI-3 NICO problem
In the NICO problem, due to the special data distribution, traditional visual deep neural networks also tend to overly focus on background information and elements when learning foreground object classification tasks. In the NICO dataset, each type of subject has its unique domain, and deep networks optimized using cross-entropy loss inherently concentrate on learning the distinctions between positive and negative examples. Therefore, it is understandable that the background, as a prominent distinction between positive and negative examples, is emphasized and learned by the deep networks.
VI-4 Summary
In summary, most end-to-end deep learning models focus more on the differences between positive and negative examples in the data, rather than the reasoning logic behind why positive examples are correct and negative examples are incorrect. When the quality of negative examples in the data is not high enough, the distinction between positive and negative examples is unable to accurately convey the underlying logic.
VII Define new negative examples to delineate the Deep-level concept (D3C)
Previous works, including Lico-Net(-Bongard), have primarily focused on designing deep learning network architectures that can achieve high reasoning accuracy on low-quality negative examples. However, this paper introduces new perspectives and a novel approach to addressing this issue.
Given that we have discovered the quality of negative examples is crucial for deep learning networks to learn the deep concepts within data, we can set more and higher-quality negative examples for our target data to enhance the performance of deep learning models. Therefore, the methodology of this section is quite straightforward: when a deep network learns from an individual observed sample in the data, we integrate other observed samples as negative samples into the current sample.
VII-A RPM and Bongard-Logo
When addressing the RPM problem and the Bongard-Logo problem, for a given observed sample, we will select images from two other distinct observed samples and incorporate them into the option pool of the original sample. The detailed procedure is illustrated in Figure 10.
Figure 10: The structure of Lico-Net-Bongard
In Figure 9, (a) illustrates the approach for handling RPM problems, while (b) depicts the method for Bongard-Logo problems. Due to computational constraints, this paper opts to augment the negative examples with three additional samples for RPM and two additional samples for Bongard-Logo.
VII-B NICO
For NICO, we cannot introduce images from outside the original dataset as negative samples, as this would disrupt the distribution of the original NICO data and undermine its significance. This is also the reason why we adhere to not importing pre-trained parameters for the deep learning model. However, our approach of introducing higher-quality negative samples remains unchanged.
When processing NICO data with a basic deep learning model, the comparison between positive and negative samples is implicitly achieved through cross-entropy. Therefore, this paper proposes that, when deep learning models process NICO data, we should introduce an explicit process of positive and negative sample comparison for the models. This is a way to improve the quality of negative samples without introducing additional data or extra supervisory signals. The explicit positive and negative sample comparison method proposed in this paper is the The Group-based Step-wise Comparison Learning Method (GSCLM). The design of GSCLM is quite straightforward. It separates the training of the perceptual part (convolutional layers) and the inductive part (fully connected layers) of the visual deep learning model, such as Vgg16.
VII-B1 The training of the perceptual part
After the perceptual part of a deep learning network processes a NICO image $x_{\text{ori}}$ into its representation $z_{\text{ori}}$, this paper selects a representation $z_{\text{pos}}$ of a same-category image $x_{\text{pos}}$ and representations $\{z_{\text{neg}_{m}}\}_{m=1}^{M}$ of multiple different-category images $\{x_{\text{neg}_{m}}\}_{m=1}^{M}$ to form a representation group. In Figure 11, we provide an example of a representation group.
Figure 11: An instance of a representation set.
In the figure, “$\phi$” represents the perceptual part of a visual deep learning network.
Subsequently, we impose loss constraints on this representation group to optimize the perceptual component of the deep network. The loss constraints can be expressed as:
${\ell_{\mathbf{D3C}}}({z_{pos}},{z_{ori}},\{{z_{ne{g_{m}}}}\}_{m=1}^{M})$
$=-\log\frac{{{e^{({z_{pos}}\cdot{z_{ori}})/\tau}}}}{{{e^{({z_{pos}}\cdot{z_{ori}})/\tau}}+\sum\nolimits_{m=1}^{M}{{e^{({z_{pos}}\cdot{z_{ne{g_{m}}}})/\tau}}}}}$
$-\log\frac{{{e^{({z_{pos}}\cdot{z_{ori}})/\tau}}}}{{{e^{({z_{pos}}\cdot{z_{ori}})/\tau}}+\sum\nolimits_{m=1}^{M}{{e^{({z_{ori}}\cdot{z_{ne{g_{m}}}})/\tau}}}}}$
$+{\ell_{{\rm{cov}}}}\left(\{{z_{pos}}\cup{z_{ori}}\cup\{{z_{ne{g_{m}}}}\}_{m=1}^{M}\}\right)$ (3)
Where $\tau$ is a temperature coefficient with a value of 0.01, $M$ represents the number of negative samples in the representation group, which is set to 9 in this paper, and ${\ell_{{\rm{cov}}}}$ is a common loss constraint term for the correlation between dimensions within the representation, which is utilized in VICReg[42].
The design of these loss constraints may resemble traditional contrastive learning algorithms; however, it is crucial to highlight that, in contrast to the emphasis on a large number of negative samples in SimCLR and the large batch learning in ViCReg, the representation group contrast approach proposed in this paper requires only a minimal number of negative samples for participation.
VII-B2 The training of the inductive part
After the perceptual component of the deep network is well-trained, we freeze these parameters and subsequently begin training the inductive part of the deep network.
VIII Designing More Demanding Negative Examples to Drive Deep-Level Concept Learning (D4C)
The methodology presented in the D3C section for RPM problems and Bongard-Logo problems is undoubtedly rudimentary. Compared to the methodology for the NICO problem, its effectiveness in improving the quality of negative examples remains uncertain. Whether and to what extent the new negative examples introduced by D3C for RPM and Bongard-Logo problems advance the development of deep concepts is still uncertain. Therefore, this paper proposes a more reasonable approach.
We have analyzed that the reason why data fails to convey precise concepts for deep models to learn is the poor quality of negative examples. So how can we obtain higher-quality negative examples without introducing additional new data or more supervisory signals? We have emphasized the importance of the quality of negative examples multiple times, and in fact, the metric reflecting the quality of an example is the probability feedback provided by the inference engine. Therefore, this paper proposes that we can design an additional sample generator to produce negative examples that receive higher evaluations from the inference engine compared to the original negative examples. The framework that integrates the generator with the inference engine can be illustrated by Figure 12.
Figure 12: The framework that integrates the generator with the inference engine
As can be seen from the figure, we have adopted the traditional framework of adversarial generation but with key modifications. By defining the structure of the generator, we can present our methodology.
It is important to note that negative examples are not unique, and consequently, our negative example generator must function as a one-to-many mapping. This poses a challenge for traditional deep learning structures to effectively perform this task, as one-to-many mappings are non-differentiable and cannot be optimized using conventional gradient descent algorithms. Therefore, this paper designs a representation generator as shown in the Figure 13.
Figure 13: The figure illustrates one block of the negative example representation generator designed in this paper.
Figure 13, taking RPM problems as an example, demonstrates the basic block structure of the negative example generator designed in this paper. It is evident that this structure is also applicable to generating negative examples for Bongard-Logo and NICO problems.
As illustrated in the Figure 13, our negative example generator does not directly generate images; instead, it generates representations of negative examples. The figure shows that after combining the representations of all images in an RPM problem matrix with a promoter and mapping them into query, key, and value vectors ($Q_{n},K_{n},V_{n}$), we use a single-layer Transformer-Encoder to compute the logical connections between them. Subsequently, within the logical output of the Transformer-Encoder, the representation corresponding to the promoter ($Q^{\prime}_{n}$) is utilized as the query to compute the step-wise linear attention mechanism with the key-value pairs formed by the representations of all images in other RPM problem matrices ($K_{n-1},V_{n-1}$). The computation process of Step-wise linear attention mechanism can be represented as:
${\text{O}_{s}}$ $=\frac{\sigma(Q^{\prime}_{n})M_{n-1}}{\sigma(Q^{\prime}_{n})z_{n-1}}$ (4)
$M_{n}$ $=M_{n-1}+\sigma(K_{n})^{T}\left(V_{n}-\frac{\sigma(K_{n})M_{n-1}}{\sigma(K_{n})z_{n-1}}\right)$ (5)
$z_{n}$ $=z_{n-1}+\sum\sigma(K_{n})$ (6)
Formulas (5) and (6) implicitly involve the participation of $k_{n-1}$ and $V_{n-1}$. This paper use element-wise ELU + 1 as the activation function $\sigma(\cdot)$. After the step-wise linear attention mechanism calculation is completed, the resulting $O_{s}$ is processed by a Feedforward layer, which follows the classic design from the Transformer, and then summed with $Q^{\prime}_{n}$ to obtain the updated promoter. The updated promoter can serve as the representation of the generated negative example, or it can also act as the input promoter for the block stacked after the current block.
Our architecture adopts a pattern similar to the infinite attention mechanism and incorporates an additional memory bank based on the linear attention mechanism to store the representations of images from the training set. As training iterations proceed, the step-wise linear attention mechanism continuously stores the image representations read during previous training batches. It is worth noting that during the first training batch, we do not compute $O_{1}$, but we still update $M_{1}$ and $z_{1}$. The update process can be explained as follows:
$M_{1}$ $=\sigma(K_{1})^{T}V_{1}$ (7)
$z_{1}$ $=\sum\sigma(K_{1})$ (8)
After stacking the generator blocks depicted in Figure 13 $S$ times without sharing parameters, a complete generator is formed. An initial prompter, ${Prompter}_{0}$, sampled from a normal distribution with optimizable mean and variance, is progressively updated through each of the $S$ blocks in the generator until it evolves into ${Prompter}_{S}$. This final ${Prompter}_{S}$ is utilized as the generated negative representation. It is not difficult to see that the generator designed in this paper has an advantage: variations in the initial promptor can lead to diversity in the generated negative examples.
The stacked negative example generator is trained in tandem with our inference engine, Lico-Net or Lico-Net-Bongard, to improve their inference accuracy. During training, the inference engine and the negative example generator are trained alternately, which is consistent with the standard adversarial learning approach. It is noteworthy that during the generator’s training, we aspire for the generated negative examples to score higher in the inference engine than the original negative examples. Therefore, we utilize cross-entropy loss to optimize the generator. Based on the principle that unsupervised learning does not ensure disentanglement [45], the design of such a loss function indicates that this paper is not concerned with the specific attributes or concepts conveyed by the generated negative examples. Instead, it leverages the fuzziness of these negative examples to push the boundaries of concepts, which constitutes a contribution to the methodology of fuzzy systems.
IX Experiment
All experiments conducted in this paper were programmed in Python, utilizing the PyTorch[46] framework.
IX-A Experiment on NICO
This paper conducted experiments on NICO, which included exploring the performance changes of ResNet18, ResNet50, VGG16, and VGG19 when combined with the D3C and D4C algorithms.
D3C enhances the positive-negative example contrast process in traditional visual deep learning models and thereby improves the quality of negative examples. Therefore, to attach D4C to these visual models, D3C must serve as the foundation. When using the D4C method to provide high-quality negative examples for the visual model processing NICO data, we maintain a memory bank for each category in the NICO dataset, that is, we maintain an $M_{n-1}$ and $Z_{n-1}$ for each category. When generating high-quality negative example representations for a specific subject category in NICO, we input the sample representations of other categories into the negative example generator. Then, the corresponding memory banks $M_{n-1}$ and $Z_{n-1}$ are selected based on the subject category to assist in the generation process. The experimental results are recorded in Table IV.
TABLE IV: D3C and D4C on NICO 8:2
dataset backbone Backbone(%) Backbone +D3C(%) Backbone +D3C+D4C(%)
NICO-Animal ResNet18 28.56$\pm$0.30 36.79$\pm$0.10 37.50$\pm$0.05
ResNet50 39.71$\pm$0.30 36.43$\pm$0.07 37.30$\pm$0.10
Vgg16 48.66$\pm$0.20 63.38$\pm$0.04 64.45$\pm$0.08
Vgg19 46.11$\pm$0.09 61.45$\pm$0.06 62.85$\pm$0.08
NICO-Vehicle ResNet18 42.29$\pm$0.08 47.82$\pm$0.05 51.91$\pm$0.06
ResNet50 55.16$\pm$0.35 47.82$\pm$0.10 48.90$\pm$0.10
Vgg16 58.24$\pm$0.40 68.09$\pm$0.14 69.75$\pm$0.15
Vgg19 57.60$\pm$0.80 70.42$\pm$0.06 71.40$\pm$0.12
  • 8 source domains for training and 2 target domains for testing
Previous works have utilized saliency maps to demonstrate the ability of their designed deep learning algorithms to identify foreground objects in NICO data. In this paper, we present saliency maps generated by the VGG16 network enhanced with D3C and D4C method. The saliency maps are shown in Figures 14 and 15.
Figure 14: The figure shows the saliency map of NICO-Animal sample.
Figure 15: The figure shows the saliency map of NICO-Vehicle sample.
This paper also compares our method with a series of approaches that do not rely on domain labels. Despite these methods incorporating pre-trained parameters into their backbones, our D3C and D4C methods outperform them when applied to the VGG16 network without the use of pre-trained parameters. The comparison results are recorded in Table V.
TABLE V: various net on NICO 8:2
data set model average accuracy
NICO-Animal and NICO-Vehicle JiGen 54.72
M-ADA 40.78
DG-MMLD 47.18
RSC 57.59
ResNet-18(pre-training) 51.71
StableNet 59.76
Vgg16+D3c+D4C 67.10
Vgg19+D3c+D4C 67.12
  • 8 source domains for training and 2 target domains for testing
IX-B Experiment on RAVEN
We tested Lico-Net on the RAVEN and I-RAVEN datasets to showcase its potential. In our exploration of the maximum potential of Lico-Net in solving RPM problems, we decided to discard the use of Sinkhorn distance for de-fuzzification and instead adopted a Transformer-Encoder equipped with learnable class vectors. Initially, we employed the Sinkhorn distance in preliminary experiments to investigate how Lico-Net induces RPM concepts. However, in order to accelerate model convergence and enhance its flexibility, we have chosen the learnable module of Transformer-Encoder to undertake the task of de-fuzzification.
Additionally, we conducted experiments combining D3C and D4C with Lico-Net to showcase the effectiveness of the D3C and D4C methods. For a fair comparison, we used the same settings and equipment as the current state-of-the-art (SOTA) method, Triple-CFN[37], including data volume, optimizer parameters, data augmentation, batch size, and other hyperparameters.
TABLE VI: Reasoning Accuracies on RAVEN and I-RAVEN.
Test Accuracy(%)
Model Average Center 2 $\times$ 2 Grid 3 $\times$ 3 Grid L-R U-D O-IC O-IG
SAVIR-T [27] 94.0/98.1 97.8/99.5 94.7/98.1 83.8/93.8 97.8/99.6 98.2/99.1 97.6/99.5 88.0/97.2
SCL [26, 27] 91.6/95.0 98.1/99.0 91.0/96.2 82.5/89.5 96.8/97.9 96.5/97.1 96.0/97.6 80.1/87.7
MRNet [24] 96.6/- -/- -/- -/- -/- -/- -/- -/-
RS-TRAN[31] 98.4/98.7 99.8/100.0 99.7/99.3 95.4/96.7 99.2/100.0 99.4/99.7 99.9/99.9 95.4/95.4
Triple-CFN[37] 98.9/99.1 100.0/100.0 99.7/99.8 96.2/97.5 99.8/99.9 99.8/99.9 99.9/99.9 97.0/97.3
Triple-CFN+Re-space[37] 99.4/99.6 100.0/100.0 99.7/99.8 98.0/99.1 99.9/100.0 99.9/100.0 99.9/99.9 98.5/99.0
Lico-Net 98.7/99.1 100.0/100.0 99.5/99.6 95.9/97.0 99.8/99.9 99.8/99.9 99.9/99.9 96.2/97.4
Lico-Net+D3C 98.9/99.3 100.0/100.0 99.7/99.8 96.2/97.5 99.9/100.0 99.9/100.0 99.9/99.9 96.8/98.0
Lico-Net+D4C 99.5/99.7 100.0/100.0 99.7/99.8 98.2/99.2 99.9/100.0 99.9/100.0 99.9/99.9 98.6/99.2
An interesting observation from the Table VI indicates a commonality among existing excellent RPM inference engines, including Lico-Net. Specifically, these inference engines exhibit varying reasoning accuracy when tackling the RAVEN [17] and I-RAVEN [40] datasets. The I-RAVEN dataset was designed to address certain limitations in the option pool configuration of RAVEN by introducing improvements to its negative example setup. These enhancements include, but are not limited to, increasing the number of attributes that are altered in the negative examples and augmenting the magnitude of attribute deviations within them. However, from a human perspective, since the setup of positive examples remains unchanged, such additional setups do not alter the difficulty of reasoning, making I-RAVEN neither more nor less challenging than RAVEN. This paper speculates that the disparate outcomes achieved by these models on RAVEN and I-RAVEN could potentially be attributed to the higher quality of negative examples present in the I-RAVEN dataset compared to those in RAVEN. This phenomenon, to some extent, also supports the analysis presented in this paper.
The Table VI also conveys that Lico-Net demonstrates impressive performance as an Inference engine. Furthermore, D4C not only enhances Lico-Net’s performance but also, to some extent, bridges the gap in Lico-Net’s performance between the RAVEN and I-RAVEN datasets.
IX-C Experiment on Bongard-Logo
We conducted experiments using Lico-Net-Bongard on the Bongard-Logo dataset; additionally, we explored the enhancement effects of the D3C and D4C algorithms on a series of excellent Bongard-Logo reasoning engines, including Lico-Net-Bongard. The experimental results are presented in the Table VII.
TABLE VII: Reasoning Accuracies of models on Bongard-logo.
Accuracy(%)
Model Train FF BA CM NV
SNAIL 59.2 56.3 60.2 60.1 61.3
ProtoNet 73.3 64.6 72.4 62.4 65.4
MetaOptNet 75.9 60.3 71.6 65.9 67.5
ANIL 69.7 56.6 59.0 59.6 61.0
Meta-Baseline-SC 75.4 66.3 73.3 63.5 63.9
Meta-Baseline-MoCo[44] 81.2 65.9 72.2 63.9 64.7
WReN-Bongard 78.7 50.1 50.9 53.8 54.3
SBSD 83.7 75.2 91.5 71.0 74.1
PMoC 92.0 90.6 97.7 77.3 76.0
Triple-CFN 93.2 92.0 98.2 78.0 78.1
Lico-Net-Bongard 92.3 91.0 98.3 76.8 77.0
Lico-Net-Bongard+D3C 92.9 91.3 98.4 77.2 77.3
Lico-Net-Bongard+D4C 93.2 92.0 98.8 78.5 77.9
Triple-CFN+D3C 93.9 92.5 98.5 78.2 78.5
Triple-CFN+D4C 94.3 92.8 98.8 78.7 79.0
PMoC+D3C 92.5 91.2 98.0 77.7 76.7
PMoC+D4C 93.0 92.6 98.2 78.0 77.5
It is noteworthy that, through experiments, we found that the negative example generator proposed in the D4C algorithm excels at generating fuzzy representations of high-scoring negative examples. This often leads to the generator easily causing gradient explosion during the training of Triple-CFN and our designed Lico-Net-Bongard. The solution presented in this paper is to perform a weighted summation of the generated negative example representations and the representations of the negative examples provided in the dataset, with a weight ratio of 0.2 to 0.8.
Finally, this paper notes that augmentation experiments were conducted for both PMoC [36] and Triple-CFN [37]. These experiments involved expanding the data volume of the Bongard-Logo dataset using data augmentation techniques, specifically including rotation and flipping. It should be emphasized that these augmentation techniques do not explicitly alter the diversity of concepts or the proportion between concepts inherent in the Bongard-Logo dataset [36]. Rather, the aim of these experiments was to investigate the potential of the models with a larger dataset. To verify the effectiveness of D4C on the augmented dataset, we conducted experiments, and the results are presented in Table VIII.
TABLE VIII: Reasoning Accuracies of D4C on Augmented Bongard-logo.
Accuracy(%)
Model Train FF BA CM NV
PMoC 94.5 92.6 98.0 78.3 76.5
Triple-CFN 94.9 93.0 99.2 80.8 79.1
Lico-Net-Bongard 95.3 94.1 99.3 81.5 78.8
PMoC+D4C 96.0 95.2 99.2 84.8 82.8
Triple-CFN+D4C 97.7 94.8 99.6 82.8 83.1
Lico-Net-Bongard+D4C 97.9 95.0 99.6 84.1 84.0
Surprisingly, Table VIII demonstrates that Lico-Net-Bongard exhibits superior performance on the augmented Bongard-Logo problem, suggesting that Lico-Net-Bongard, designed based on fuzzy system methodology [14, 15], may possess greater network capacity and stronger learning abilities. Especially when combined with the D4C algorithm, Lico-Net-Bongard’s performance on the augmented Bongard-Logo problem is particularly outstanding, fully proving the effectiveness and potential of fuzzy systems as a methodological guide for deep learning algorithms.
IX-D Experiment on PGM
In this paper, we conducted experiments related to Lico-Net on PGM. And the results are recorded in Tabel IX. The data recorded in the table reveals many meaningful phenomena.
TABLE IX: Reasoning Accuracies of models on PGM.
Model Test Accuracy(%)
SAVIR-T [27] 91.2
SCL [26, 27] 88.9
MRNet [24] 94.5
RS-CNN[31] 82.8
RS-TRAN[31] 97.5
Triple-CFN[37] 97.8
Triple-CFN+Re-space layer[37] 98.2
RS-TRAN+D3C 97.8
RS-TRAN+D4C 98.3
Triple-CFN+Re-space layer+D3C 98.3
Triple-CFN+Re-space layer+D4C 98.8
Lico-Net 97.9
Lico-Net+D3C 98.2
Lico-Net+D4C 99.0
The combinations of Lico-Net+D3C, Triple-CFN+D3C, and RS-Tran+D3C exhibited very limited improvement in reasoning accuracy, which was expected. After all, there are fundamental differences between RPM and Bongard-Logo. RPM problems share a common concept space and attribute recognition mechanism, while D3C, designed to address concept interference and erroneous coupling issues between different samples, is less effective for RPM problems, as anticipated. In contrast, D4C brought significant performance improvements to Lico-Net, Triple-CFN, and RS-Tran. This fully validates our hypothesis that enhancing the quality of negative samples can effectively boost the performance of deep learning algorithms. This finding holds important theoretical and practical significance.
Additionally, we conducted experiments on the generalization issue in PGM for Lico-Net, D3C, and D4C. The experimental results are recorded in Table X.
TABLE X: Generalization Results of Lico-Net in PGM.
Accuracy(%)
Dataset Lico-Net Lico-Net+D3C Lico-Net+D4C
Interpolation 81.5 82.0 82.9
Extrapolation 18.5 18.6 18.6
Held-out Attribute shape-colour 12.8 13.5 13.6
Held-out Attribute line-type 26.0 26.5 27.5
Held-out Triples 23.0 23.5 24.3
Held-out Pairs of Triples 44.8 45.0 45.5
Held-out Attribute Pairs 29.5 30.0 30.4
IX-E Ablation experiments
The negative example generator in this paper is supported by a step-wise linear attention mechanism, and the performance improvement it brings to the generator needs to be clarified. Therefore, ablation experiments were conducted in this paper.
We removed the linear attention mechanism from the D4C generator while retaining the process of the transformer-encoder for processing input representations, and conducted relevant experiments again. The experimental results are recorded in the entry labeled “D4C-” in the Table XI, XII and XIII.
TABLE XI: Ablation studies on PGM.
Model Test Accuracy(%)
RS-TRAN+D4C 98.3
Triple-CFN+Re-space layer+D4C 98.8
Lico-Net+D4C 99.0
RS-TRAN+D4C- 97.9
Triple-CFN+Re-space layer+D4C- 98.5
Lico-Net+D4C- 98.7
TABLE XII: Ablation studies on Augmented Bongard-logo.
Accuracy(%)
Model Train FF BA CM NV
PMoC+D4C 96.0 95.2 99.2 84.8 82.8
Triple-CFN+D4C 97.7 94.8 99.6 82.8 83.1
Lico-Net-Bongard+D4C 97.9 95.0 99.6 84.1 84.0
PMoC+D4C- 95.0 93.3 98.5 79.0 77.8
Triple-CFN+D4C- 95.5 94.0 99.3 81.0 80.2
Lico-Net-Bongard+D4C- 96.0 94.2 99.3 82.0 80.8
TABLE XIII: Ablation studies on NICO 8:2
dataset backbone Backbone +D3C(%) Backbone +D3C+D4C(%) Backbone +D3C+D4C-(%)
NICO-Animal ResNet18 36.79$\pm$0.10 37.50$\pm$0.05 37.10$\pm$0.04
ResNet50 36.43$\pm$0.07 37.30$\pm$0.10 36.60$\pm$0.08
Vgg16 63.38$\pm$0.04 64.45$\pm$0.08 64.00$\pm$0.04
Vgg19 61.45$\pm$0.06 62.85$\pm$0.08 62.10$\pm$0.05
NICO-Vehicle ResNet18 47.82$\pm$0.05 51.91$\pm$0.06 49.32$\pm$0.10
ResNet50 47.82$\pm$0.10 48.90$\pm$0.10 48.10$\pm$0.10
Vgg16 68.09$\pm$0.14 69.75$\pm$0.15 68.62$\pm$0.18
Vgg19 70.42$\pm$0.06 71.40$\pm$0.12 70.60$\pm$0.20
  • 8 source domains for training and 2 target domains for testing
It can be observed from the experimental results recored in these three tables that the step-wise linear attention mechanism serves as a crucial external memory bank for the D4C negative example generator, providing stable performance improvements.
X Conclusion
This paper points out that when deep learning is applied to problems that embed complex and deep human concepts, deep learning algorithms should explicitly extract these concepts as fuzzy representations and reason based on these representations. This is also a key methodology in fuzzy systems. The Lico-Net and Lico-Net-Bongard designed in this paper both demonstrate the validity of this viewpoint and prove that fuzzy systems should be one of the key development directions for deep learning.
More notably, this paper proposes that deep learning models tasked with discriminative tasks are more concerned with the distinctions between positive and negative examples rather than the logic of discriminating between them. Focusing on the distinctions between positive and negative examples is feasible and effective in traditional discriminative tasks. However, when the quality of negative examples is not high enough to strictly define the boundaries of concepts, the performance of the algorithm may be compromised. Therefore, this paper suggests that providing stricter negative examples for deep learning models can effectively improve their performance. Consequently, this paper designs the GSCLM algorithm for NICO data and a negative example generator based on a step-wise linear attention mechanism for RPM problems, Bongard-Logo problems, and NICO problems. The results of the experiment demonstrate the effectiveness of these algorithms. Specifically, this generator represents a novel attempt within the methodology of fuzzy systems, and its effectiveness hints at the potential of fuzzy systems in the field of artificial intelligence.
Thus, this paper advocates that while pursuing the design of algorithms that can extract strong performance from low-quality negative examples, the approach of generating high-quality negative examples and enhancing the contrast between positive and negative examples proposed in this paper is also worthy of attention and consideration.